mrs. tate bought 5 1/2 of candy to divide among her 22 students. if the candy was divided equally, how many pounds of candy did each student receives?
Answer by macston(5194)

5 1/2 lbs=11/2 lbs
11/2 lbs* 1/22=11/44 lbs=1/4 lb ANSWER Each student would receive 1/4 pound of candy.